Two drones are used to deliver parcels in succession. The probability of each drone correctly delivering the parcel assigned to it is 0.3 and 0.2 respectively. Find the probability that the parcel is delivered



1
Expert's answer
2021-11-17T17:50:20-0500

Let "A" denote the event "first drone delivers the parcel correctly".

Let "B" denote the event "second drone delivers the parcel correctly".

Given "P(A)=0.3, P(B)=0.2." Since the events are independent, then


"P(A\cap B)=P(A)P(B)"

Find the probability that the parcel is delivered


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=P(A)+P(B)-P(A)P(B)"

"=0.3+0.2-0.3(0.2)=0.44"